OpenStudy (anonymous):

fan and medal !!!the product of (a-b)(a-b) is square?? always sometimes never

OpenStudy (anonymous):

product of (x-5)^2

OpenStudy (anonymous):

product of (x+5)^2

OpenStudy (anonymous):

product of (2x-5y)^2

OpenStudy (anonymous):

product of (4x+3y)^2

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@joshhfers @Johnbc

OpenStudy (anonymous):

@navk

OpenStudy (anonymous):

(X-5)^2 is the same as (x-5)(x-5) First you multiply x with (x-5) and than you multiply 5 with (x-5) and you subtract the outcome of that with the first answer. The other questions are all the same type of question.

OpenStudy (anonymous):

always

OpenStudy (anonymous):

what about the + ones that would be like (x+5)(x+5)

OpenStudy (anonymous):

if you consider a value in parenthesis multiplied by the same value in the same parenthesis then you can say it is like multiplying the same value by itself which is the same as \[x \times x = x^2\]

OpenStudy (anonymous):

Another example \[(100x+74639) \times (100x + 74639)\] Is tied together because of the parenthesis so we can say it is the same as a value multiplied by itself and we know that any value multiplied by itself will be squared so long as it is only multiplied by itself once and not a second time because then that would mean we have "cubed" of our value or\[x \times x \times x = x^3\]